OpenStudy (anonymous):

ryleec i need your help please The temperature at 8:00 a.m. was –6°F. By 4:00 p.m. the temperature was 17°F. What was the change in temperature from 8:00 a.m. to 4:00 p.m.? A. –23°F B. –11°F C. 11°F D. 23°F

OpenStudy (xapproachesinfinity):

what's the difference in time to start this

OpenStudy (xapproachesinfinity):

8am to 4 pm how many hours are there

OpenStudy (xapproachesinfinity):

okay so from 8 am to 4 pm there are 4 hours

OpenStudy (xapproachesinfinity):

I mean 8 hours* now final T=17 F initial T=-6 F the diffrence is 17-(-6)=17+6=23 F so

OpenStudy (xapproachesinfinity):

Oh i miss understood the problem all you need is the difference in temperature \(\large \rm T_f-T_i=23^\circ F\)

OpenStudy (anonymous):

it's d